In Problems 1 through 23, find an eigenvector corresponding to each eigenvalue of the given matrix.

15. 3 0 -17 23 2 -1 ( 3 2 .

5 -7 17.4 -1 -3 2

0 2010 0 10 01 -10

24. Find unit eigenvectors (i.e., eigenvectors whose magnitudes equal unity) for the matrix in Problem 1.

1. [-1 ]
Answer #1

(15)

(17)

17) er - 7 7 -3 4 4 2 is 5-1 7. 4 -3-) =(5- -1 2-2 The characteristic polynomial (7»{-(30)(2-) 693-746-460-03 +7(-5+4 (3+2) =(A-3910) ~ 1 O -1 RIR + 2 R2 - 0 R₂ + R3-7R2 X, xz - O. 70 X2 - 3 . The solution set as X3 let, z = 1, Ba for 325, B= --- is

(20)

ther, the rector, if By be eigen Bz z 2 eigenvalues Corresponding eigeneton - o i i -

(1)
Answer #1

(15)

(17)

17) er - 7 7 -3 4 4 2 is 5-1 7. 4 -3-) =(5- -1 2-2 The characteristic polynomial (7»{-(30)(2-) 693-746-460-03 +7(-5+4 (3+2) =(A-3910) ~ 1 O -1 RIR + 2 R2 - 0 R₂ + R3-7R2 X, xz - O. 70 X2 - 3 . The solution set as X3 let, z = 1, Ba for 325, B= --- is

(20)

ther, the rector, if By be eigen Bz z 2 eigenvalues Corresponding eigeneton - o i i -

(1)